OpenStudy (anonymous):

how can I solve a 2 step equation? plz include an example

OpenStudy (anonymous):

what do you mean by 2 step equation

OpenStudy (anonymous):

something like that Ax+b = Z ??

OpenStudy (anonymous):

like 3x+4=15

OpenStudy (anonymous):

so the 3x+4= 15 became 3x= 15 -4 became 3x=11 and the x=11/3 that all foolow those step on every equation like this one

OpenStudy (anonymous):

thank you